### v1

# Compact Speech Translation Models via Discrete Speech Units Pretraining

Abstract
Using Self-Supervised Learning (SSL) as model initialization is now common to obtain strong results in Speech Translation (ST). However, they also impose a large memory footprint, hindering on-device deployment. In this paper, we leverage the SSL models by pretraining smaller models on their Discrete Speech Units (DSU). We pretrain encoder-decoder models on 1) Filterbank-to-DSU and 2) DSU-to-Translation data, and take the encoder from 1) and the decoder from 2) to initialise a new model, finetuning this on limited speech-translation data. The final model becomes compact by using the DSU pretraining to distil the knowledge of the SSL model. Our method has several benefits over using DSU as model inputs, such as shorter inference pipeline and robustness over (DSU) tokenization. In contrast to ASR pretraining, it does not require transcripts, making it applicable to low-resource settings. Evaluation on CoVoST-2 X-En shows that our method is >$0.5$ BLEU better than a ST model that directly finetune the SSL model, given only half the model size, and on a par with ASR pretraining.
Compact Speech Translation Models via Discrete Speech Units Pretraining
Tsz Kin Lam and Alexandra Birch and Barry Haddow School of Informatics, University of Edinburgh {tlam, a.birch, bhaddow}@ed.ac.uk
1 Introduction
In Speech-to-Text Translation (ST), the use of SSL models, e.g., wav2vec2 and HuBERT Baevski et al. (2020); Hsu et al. (2021), as model initialization is now common to obtain SOTA result Agarwal et al. (2023). Nevertheless, using these bulky models as initialisation imposes a large memory footprint to the downstream ST, which limits flexibility in modelling choices. In addition, it hinders on-device deployment that is crucial for privacy and useful in the absence of internet connection.
Knowledge distillation (KD) Kim and Rush (2016); Inaguma et al. (2021) is the standard approach to creating smaller models, but obtaining the high-quality pseudo labels for training ST systems is very costly. Instead, we focus on Discrete Speech Units (DSU) since they can be obtained cheaper, e.g., without requiring a translation model, and are an intermediate representation between speech and text. DSUs are K-means clusters of speech representations from selected layers within a SSL model. It represents sequences of discrete tokens, which are easier to model within a text processing architecture Polyak et al. (2021); Chou et al. (2023). DSU sequences are far smaller than dense Filterbank sequences, and they can be shortened with removing duplicates and Byte Pair Encoding (BPE) Sennrich et al. (2016). Therefore, in ST, a straightforward method to distill the SSL models is to replace the dense Filterbank by the DSUs, aka the DSU-to-Translation (DSU-to-Trl) model. Although using DSUs for training allows for transfer learning, using them at inference requires large and slow models.
To address the above, we use DSU for pretraining rather than as model input to make ST models more compact. Our method leverages large SSL models by pretraining smaller models on the corresponding DSU. More specifically, it pretrains encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Trl data, and takes the encoder from 1) and the decoder from 2) to initialise a new model. After that, the new model is finetuned on limited speech-translation data. Under this formulation, (1) the model is much more compact than the SSL model (2) the compact model does not use DSU as inputs which (a) avoids the lengthy SSL and clustering inference pipeline and (b) minimizes its sensitivity to (DSU) tokenization. (3) Our method does not require transcripts, unlike ASR pretraining, making it applicable to low-resource settings. Similar to ST methods that use pretrained components, our method could be limited by the pretraining modality gap Liu et al. (2020); Le et al. (2023). Motivated by prior works, we investigate mitigating it with Connectionist Temporal Classification (CTC) Graves et al. (2006) regularisation.
We evaluate our method on CoVoST-2 Wang et al. (2021) X-En (21 language directions) with multilingual ST. By simply using a monolingual HuBERT model to extract the DSU, our method has a gain of >$3$ BLEU with respect to training from scratch, demonstrating DSU’s cross-lingual ability. Our method is also >$0.5$ BLEU better than a ST model that directly finetuned HuBERT despite having half model size. Our main contributions are:
  • We propose to distill knowledge from large SSL models to make ST models more compact via DSU pretraining. Our method avoids the lengthy inference pipeline in the DSU-to-Trl method.
  • We examine the effect of (DSU) tokenization on the ST models. Our method is robust across different tokenizations, which is not the case for the DSU-to-Trl method.
  • We examine applying CTC to the DSU in pretraining and on the translation in finetuning to mitigate the modality gap, and we show that both improve performance.
2 Related Work
There are a number of related works that use DSU to enhance ST. Fang and Feng (2023) and Zhang et al. (2023b) use DSU to create more training data in a back-translation fashion. Chang et al. (2023) and Zhang et al. (2023b) explore the replacement of Filterbank by DSU as speech input. In contrast, we use DSU and its translation model for pretraining, resulting in a better Fbk-to-Trl model that has a shorter inference pipeline.
In the case of pretraining, Wu et al. (2023) use a single Speech-to-DSU model in pretraining for general speech-to-text purposes whereas we tailor the use for ST by using a pair of encoder-decoder models. Zhang et al. (2022b) also decompose ST into speech-to-unit and unit-to-text tasks. Their training is based on masked unit prediction, and it requires an extra unit-encoder module in inference. In contrast, we resort to supervised training on the DSU in acoustic pretraining and require no extra module in inference. More importantly, our goal is to make (multilingual) ST more compact, aiming also at low-resource settings where transcripts are not easily available, rather than learning a joint semantic space for both transcripts and audios.
AST model (#Params) High Mid Low All
Scratch (52M) $19.37$ $7.91$ $0.73$ $5.99$
ASR-Pretraining (52M) $26.53$ $12.22$ $1.82$ $9.00$
Hu-Transformer (113M) $24.31$ $11.42$ $2.18$ $8.60$
DSU-Adapter (48M) $26.54$ $12.91$ $1.76$ $9.13$
Table 1: Results in BLEU on CoVoST-2 X-En. DSU-Adapter is 0.53 points better than Hu-Transformer that is $2.3$ times larger and on a par with ASR-Pretraining.
3 Method
Our method distills knowledge from the SSL representations via first DSU pretraining (PT) and then finetuning (FT) on the paired ST data.
In PT, our method uses DSU from an SSL speech model to train a pair of encoder-decoder models: 1) Fbk-to-DSU and 2) DSU-to-Trl. The Fbk-to-DSU model takes the log Mel Filterbank as the encoder input and predicts the corresponding DSU sequence. The model is trained by Maximum Likelihood Estimation (MLE). In addition, we interpolate between MLE and CTC applied to the last encoder layer:
$\mathcal{L}^{\text{Fbk-to-DSU}}=(1-\lambda_{\alpha})\mathcal{L}_{\text{MLE}}(\mathbf{U}|\mathbf{F})+\lambda_{\alpha}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{U}}|\mathbf{F})$ (1)
[…]
In FT, we use the encoder of the Fbk-to-DSU model and the decoder (and its output layer) of the DSU-to-Trl model to initialise the compact model, followed by finetuning on the paired speech-translation data using both MLE and CTC loss Gaido et al. (2021); Zhang et al. (2023a):
$\mathcal{L}^{\text{FT}}=(1-\lambda_{\beta})\mathcal{L}_{\text{MLE}}(\mathbf{Y}|\mathbf{F})+\lambda_{\beta}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{Y}}|\mathbf{F})$ (2)
[…]
Tokenization
As mentioned in section 1, we examine the tokenization effect, including using separate (source/target) vocabularies or not. We denote each DSU token by a hashtag with an index, e.g,. ’#456’. Following Ao et al. (2022), we reduce the sequence length via merging the sequential repetitions, e.g, ’#1 #1 #1 #456 #456 #23’ becomes ’#1 #456 #23’. When BPE is applied on the DSU, the whitespace characters are ignored, e.g., ’#1 #456 #23’ becomes ’#1#456#23’ which could be split into ’#1#45 6#2 3’.
Has BPE $|\mathcal{V}|$ DSU Length DSU-to-Trl Enc-Init EncDec-Init DSU-Adapter
on DSU? Length Ratio $20$M to $27$M $52$M to $70$M $46$M to $64$M $48$M to $67$M
No $1$K-$8$K $176$ $12.9$ $6.73$ $7.70$ $7.87$ $8.54$
$1$K-$16$K " $14.1$ $6.36$ $7.50$ $8.00$ $8.43$
$1$K-$32$K " $14.9$ $6.30$ $7.23$ $7.65$ $7.94$
$8$K " $12.7$ $6.88$ $7.64$ $8.05$ $8.26$
$16$K " $14.0$ $6.33$ $7.41$ $7.91$ $8.17$
$32$K " $14.9$ $6.26$ $6.66$ $7.41$ $7.68$
Yes $1$K-$8$K $221$ $16.3$ $4.52$ $8.23$ $8.44$ $8.61$
$16$K-$8$K $129$ $9.5$ $5.06$ $8.51$ $8.76$ $8.95$
$32$K-$8$K $115$ $8.5$ $4.43$ $8.67$ $9.02$ $9.13$
$8$K $150$ $7.6$ $7.02$ $8.33$ $8.51$ $8.82$
$16$K $133$ $7.7$ $6.50$ $8.57$ $8.61$ $8.93$
$32$K $118$ $7.8$ $5.07$ $8.30$ $8.44$ $8.70$
Table 2: Tokenization effect on 4 types of ST methods. In general, darker (brighter) cells refer to weaker (stronger) models. The best two models apply both BPE on the DSU and separate vocabularies in PT (cells in yellow).
4 Experiments
4.1 Data Preprocessing
We follow standard practices to preprocess the CoVoST-2 X-En data, i.e., log-Mel Filterbank of $80$ dimensions for speech inputs and BPE with $8$K vocabulary on the English target, $16$K on the transcripts and $32$K on the DSU, unless otherwise specified. We use an English HuBERT-Base model to extract the DSU by first downsampling the CoVoST-2 audio to $16$KHz. Each audio data utterance is then converted into the DSU, i.e., the clustering indexes, by applying K-Means clustering (K=$1,000$) on its HuBERT representation from the 6th layer Lakhotia et al. (2021). We sample training data from each language pair to form a subset of $98$K instances to train the K-Means model. Further details are in Appendix A.1.
4.2 Model Configuration
All models are based on Transformer Vaswani et al. (2017) with implementations from fairseq Ott et al. (2019); Wang et al. (2020). In the Fbk-to-token [^1] Token refers to either transcriptions, DSU, or translations. models, the encoder has convolutional layers to downsample the Fbk by a factor of 4. There are 12-6 layers in the transformer encoder-decoder, whereas the embedding and feed-forward network (FFN) dimensions are 256 and 4,096 respectively, unless otherwise specified. It is worth noting that: (1) The Fbk-to-DSU model is not trained on the translations, so it is not directly comparable to the ST models. Its effect to ST lies on its pretrained encoder (Table 2). (2) The DSU-to-Trl model is a ST model, and it can also be used for initialization.
Scratch is a ST model trained on the paired speech-translation data without pretraining.
ASR Pretraining refers to a ST model whose encoder is initialized by a speech recognition task with CTC regularisation on the transcripts.
DSU-to-Trl follows the Transformer used in text translation. We use 6-6 layers in the encoder-decoder which the dimension of embedding and FFN is 256 and 2,048 respectively. Despite its smaller model size, its inference requires storing and calling both HuBERT and K-Means models.
Hu-Transformer uses the entire HuBERT as the speech encoder initialization Fang and Feng (2023). For comparison to our DSU-Adapter, its subsequent encoder-decoder also has 1-6 layers.
DSU-Adapter is our proposed method. To better align the two pre-trained components, we also experiment with adding an extra encoder layer [^2] All parameters are trainable in ST finetuning. after the pre-trained encoder. Since its decoder is initialized by the DSU-to-Trl method, its decoder FFN dimension is $2,048$.
Enc-Init is a ST model that has its encoder initialized by the Fbk-to-DSU encoder. EncDec-Init is a DSU-Adapter model without the adapter layer.
4.3 Training and Inference
We leave the details in Appendix A.3. It is worth noting that we do not use extra audio data, e.g., Libri-Light Kahn et al. (2020) in pretraining.
5 Results and Analysis
We divide the 21 language pairs by resource level into: 1) "High": {ca, de, es, fr}, 2) "Mid": {fa, it, pt, ru and zh}, 3) "Low": {ar, cy, et, id, ja, lv, mn, nl, sl, sv, ta, tr} and 4)"All": the 21 languages pairs. We report the average BLEU over the test sets of each group using sacrebleu Post (2018)333nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1.
Table 1 compares our DSU-Adapter and the baselines. Our DSU-Adapter is $3$ BLEU points (in the group "All") higher than Scratch. Furthermore, it is better than Hu-Transformer in spite of having half the parameters. For "Mid" and "High", the improvement is $1.49$ and $2.23$ points respectively, but it falls short by $0.42$ points for "Low". We also compare to ASR pretraining, which is not always applicable, e.g., in low-resource setting or perhaps even in an unwritten language Zhang et al. (2022a). Surprisingly, our adapter is on a par with it, i.e., $0.13$ points better. The result remains consistent when measured in chrF Popović (2015) and COMET [^4] WMT22-COMET-DA Rei et al. (2022), see Appendix A.4.
5.1 Tokenization and its Transfer
Table 2 presents the effect of tokenization. The 1st column indicates if BPE is applied on the DSU, and the 2nd column "$|\mathcal{V}|$" shows the vocabulary size, and if the model has a joint vocabulary or not in PT. For example, "$1$K-$8$K" means that we use a vocabulary of size $1$K for DSU and $8$K for English. These results are for all languages, please see Appendix A.4 for fine grained results.
Effect of BPE on DSU-to-Trl
When BPE is not applied on the DSU, the DSU-to-Trl models have a BLEU of $6.48$$\pm$ $0.26$ across the 6 vocabulary settings. Despite being smaller (<$30$M), these models are better than the Scratch model of $5.99$ BLEU.
When BPE is applied, the DSU sequence length becomes a crucial factor. On one hand, it delivers the best model ("$8$K": $7.02$ BLEU). On the other hand, the result deteriorates a lot in the other 5 settings, i.e., a BLEU of $5.12$$\pm$ $0.83$. More importantly, BPE’s correlation with the model performance is not straightforward, neither is the length ratio. For an example, the "$32$K" model (DSU length of $118$) is about $2$ BLEU points behind to the "$8$K" model (DSU length of $150$).
DSU-Adapter is more tokenization robust
When BPE is applied on the DSU, DSU-Adapter has a higher BLEU of $8.86$$\pm$ $0.19$, in contrast to $8.17$$\pm$ $0.32$. This behaviour is opposed to the DSU-to-Trl model that has a lower BLEU of $5.54$$\pm$ $1.07$ (with also larger variance), in contrast to $6.48$$\pm$ $0.26$ when BPE is not used. The better scores with smaller variance show that DSU-Adapter is more tokenization robust. We see this as a benefit of introducing the DSU (SSL knowledge) via PT rather than as model inputs. Besides, using separate vocabularies in PT is preferred since it performs slightly better, and the DSU, which are not needed in the ST output, would not occupy the target vocabulary.
Ablation
Having similar size, DSU-Adapter is better than EncDec-Init and Enc-Init. The BLEU over their 12 settings is $8.51$$\pm$ $0.44$, $8.22$$\pm$ $0.51$, and $7.99$$\pm$ $0.61$ respectively. Encoder-initialization seems more crucial than decoder-initialization, as reflected by the fact that the best DSU-Adapter comes from a combination with the weakest DSU-to-Trl model of $4.43$ BLEU.
5.2 CTC regularisation helps
Has CTC in High Mid Low All
DSU PT? ST FT?
No No $25.81$ $9.91$ $1.46$ $8.14$
No Yes $26.10$ $11.35$ $1.69$ $8.71$
Yes No $25.94$ $10.82$ $1.51$ $8.44$
Yes Yes $26.12$ $11.53$ $1.73$ $8.74$
Table 3: Effect of CTC on Fbk-to-DSU PT and/or ST FT to the DSU-Adapter method.
Table 3555Owing to the large number of configurations, we can experiment on 4 vocabularies only and then report their average. presents an analysis of using CTC in the Fbk-to-DSU PT and/or in the ST FT using the DSU-Adapter method. We find that CTC helps on either stage, but the gain is 0.27 BLEU more in finetuning. Using them jointly still helps, but the gain becomes marginal, i.e, 0.03 BLEU.
6 Conclusion
We propose to distill knowledge from the large SSL model via pretraining on its DSU to make ST models more compact for on-device deployment. Our compact model, i.e., the DSU-Adapter, shows stronger result over finetuning the corresponding (larger) HuBERT in multilingual ST task. In contrast to using DSU as model input, our method requires less components in inference and shows stronger robustness over the tokenization. Finally, our method is suitable for the low-resource setting since it requires no transcripts.
7 Limitations
In this work, we propose to distill knowledge from the informative SSL representations via DSU pretraining to make ST models more compact. In addition, we also discuss advantages of our method over using DSU as model inputs. In spite of its noticeable benefits, there are several factors that are not thoroughly explored:
K-Means Clustering
We did not inspect the clustering size (fixed to $1,000$) and the number of training instances (only fixed to $98,000$) used in training the K-Means clustering model.
Other acoustic encoders
We did not experiment other acoustic encoders, such as conformer Gulati et al. (2020); Papi et al. (2023) and E-Branchformer Peng et al. (2023). This stronger encoders should provide further gains for our method since they also enjoy the benefit of pretraining.
Self-Supervised Models
We did not experiment using other layers (only fixed to the 6th layer) from the HuBERT model nor from stronger SSL models to extract the DSU. These two factors potentially increase the DSU quality and potentially provide better improvement for our method while maintaining the same model size.
Acknowledgements
This work was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10039436: UTTER). The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
Appendix A Appendix
A.1 Data preprocessing
log Mel Filterbank
The 80-D log Mel Filterbank is computed every 10ms with a 25ms window, followed by feature normalisation using statistics, i.e., mean and variance, computed over each channel.
BPE
We use the BPE implementation from sentencepiece Kudo and Richardson (2018) and obtain vocabulary of size 8K on the English target, 16K on the transcripts and 32K on the DSU, unless otherwise specified.
HuBERT
K-Means Clustering
We use MiniBatchKMeans from sklearn with a clustering size of $1,000$ and default settings in other hyper-parameters. When training the clustering model, we divide the 21 language pairs into three groups: 1) {ar, cy, et, id, ja, lv, mn, sl, sv, ta, tr}, 2) {nl, pt, ru, zh} and 3) {ca, de, es, fa, fr, it} for sampling. For each language pair in group 1), we sample $1$K instances, whereas it becomes $3$K in group 2) and $12.5$K in group 3), resulting in a total of $98$K instances.
A.2 Model Configuration
We use "pre" layer-normalization Nguyen and Salazar (2019) in DSU-to-Trl.
A.3 Training and Inference
In training,
  • we skip training data that are more than $3$K frames for Filterbank, $320$K frames for HuBERT input or $1,024$ tokens.
  • we apply SpecAugment Park et al. (2019) with parameters: {$F=30,T=40,m_{F}=2,m_{T}=2$} on Filterbank inputs.
  • we share the embedding weights when using a joint vocabulary in the DSU-to-Trl model.
  • we set $\lambda_{\alpha}$ and $\lambda_{\beta}$ in CTC to $0.3$ and the smoothing parameter in LS to $0.1$
  • we initialize the encoder (decoder) with the last (best) checkpoint from the PT model.
All models are optimized by Adam Kingma and Ba (2015) with inverse square root scheduler. For simplicity, we provide major training parameters for each model in the format of {effective batch size, warm-up steps, peak learning rate, maximum training steps}:
Fbk-to-Token:
{32K frames, 25K, 2e-3, 60K}
DSU-to-Trl:
{80K tokens, 10K, 5e-4, 50K}
Hu-Transformer:
{4M frames, 4K, 1e-4, 300K}
In inference, we average the last 5 checkpoints and use beam size of 5 in generation. All experiments are run on Nvidia A100 GPUs. It takes about 1 day for 2 A100 (40GB) GPUs to complete an experiment that uses Filterbank as inputs.
A.4 Results and Analysis
In addition to BLEU (see Table 8 for the result of each language), we also provide the result of Table 1 using chrF777nrefs:1—case:mixed—eff:yes—nc:6—nw:0—space:no—version:2.3.1 and COMET. In order to compute the COMET score, we feed the gold-reference transcripts as the source sentence.
Table 4 and Table 6 show the average result in chrF and COMET respectively over the four groups defined in section 5. Their results follow the findings in BLEU: the DSU-Adapter is better than Hu-Transformer and on a par with ASR pretraining. Further details, such as training data size (in terms of the number of utterances rather than hours888”zh” has 10 hours of audio data, whereas ”nl” has only 7 hours. Therefore, ”zh” is classified into ”Mid” resource group even though it has slightly less number of utterances than ”nl”.) and per language performance, are shown on Table 5 for chrF and Table 7 for COMET.
CTC Regularisation
On Table 3, we average the results over this four configurations: 1) "No-BPE 1K-8K", 2) "BPE 8K", 3) "BPE 32K" and 4) "BPE 32K-8K".
AST model Transcripts Model size High Mid Low All
Scratch No 52M $43.58$ $27.22$ $14.62$ $23.10$
ASR Pretraining Yes 52M $51.90$ $32.82$ $16.38$ $27.10$
Hu-Transformer No 113M $49.88$ $31.94$ $16.96$ $26.80$
DSU-Adapter No 48M $52.08$ $33.94$ $16.46$ $27.40$
Table 4: Results of Table 1 in chrF score by resource group. Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $42.5$ $50.9$ $49.8$ $51.5$
de 127,824 $35.4$ $45.1$ $43.6$ $45.4$
es 79,013 $48.1$ $56$ $52.8$ $55.7$
fr 207,372 $48.3$ $55.6$ $53.3$ $55.7$
fa 53,949 Mid $15.6$ $16.9$ $19.9$ $16.9$
it 31,698 $40.6$ $50.2$ $46.4$ $50.3$
pt 9,158 $25$ $31.3$ $33.8$ $32.3$
ru 12,112 $35.6$ $42.6$ $40$ $47.7$
zh 7,085 $19.3$ $23.1$ $19.6$ $22.5$
ar 2,283 Low $12.5$ $14.4$ $13.8$ $13.2$
cy 1,241 $14.5$ $16.3$ $15.8$ $15.4$
et 1,782 $17.9$ $21.3$ $19$ $19.9$
id 1,243 $13.8$ $14.2$ $14.9$ $14.5$
ja 1,119 $14$ $14$ $13.7$ $14.4$
lv 2,337 $13.9$ $14.5$ $15.9$ $15.4$
mn 2,067 $14.5$ $15.4$ $14.3$ $14.8$
nl 7,108 $17.8$ $24.3$ $30.8$ $27$
sl 1,843 $13.9$ $15.4$ $15.3$ $15.4$
sv 2,160 $13.2$ $13.8$ $17.3$ $14.5$
ta 1,358 $12$ $11.9$ $12$ $12.3$
tr 3,966 $17.4$ $21.1$ $20.7$ $20.7$
Table 5: Results of Table 1 in chrF score for each language pair. The best result in each group is denoted by ’_’.
AST model Transcripts Model size High Mid Low All
Scratch No 52M $0.605$ $0.498$ $0.433$ $0.481$
ASR Pretraining Yes 52M $0.680$ $0.537$ $0.443$ $0.511$
Hu-Transformer No 113M $0.650$ $0.522$ $0.439$ $0.499$
DSU-Adapter No 48M $0.681$ $0.548$ $0.442$ $0.513$
Table 6: Results of Table 1 in COMET score by resource group. Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $0.607$ $0.681$ $0.654$ $0.684$
de 127,824 $0.535$ $0.608$ $0.588$ $0.613$
es 79,013 $0.620$ $0.696$ $0.654$ $0.692$
fr 207,372 $0.660$ $0.736$ $0.704$ $0.735$
fa 53,949 Mid $0.479$ $0.494$ $0.501$ $0.493$
it 31,698 $0.557$ $0.640$ $0.594$ $0.641$
pt 9,158 $0.468$ $0.497$ $0.510$ $0.508$
ru 12,112 $0.551$ $0.580$ $0.563$ $0.630$
zh 7,085 $0.434$ $0.472$ $0.440$ $0.467$
ar 2,283 Low $0.438$ $0.465$ $0.45$ $0.445$
cy 1,241 $0.420$ $0.448$ $0.424$ $0.426$
et 1,782 $0.413$ $0.432$ $0.407$ $0.423$
id 1,243 $0.420$ $0.406$ $0.416$ $0.412$
ja 1,119 $0.417$ $0.412$ $0.411$ $0.422$
lv 2,337 $0.457$ $0.453$ $0.465$ $0.461$
mn 2,067 $0.429$ $0.441$ $0.419$ $0.437$
nl 7,108 $0.427$ $0.460$ $0.483$ 0.477
sl 1,843 $0.430$ $0.428$ $0.433$ $0.428$
sv 2,160 $0.454$ $0.448$ $0.472$ $0.454$
ta 1,358 $0.413$ $0.406$ $0.403$ $0.411$
tr 3,966 $0.480$ $0.519$ $0.484$ $0.514$
Table 7: Results of Table 1 in COMET score for each language pair. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $18.24$ $25.66$ $24.24$ $25.98$
de 127,824 $13.97$ $21.06$ $19.28$ $21.32$
es 79,013 $22.37$ $29.8$ $26.64$ $29.34$
fr 207,372 $22.91$ $29.6$ $27.09$ $29.52$
fa 53,949 Mid $1.9$ $3.07$ $4.63$ $2.84$
it 31,698 $15.1$ $23.55$ $19.98$ $23.36$
pt 9,158 $7.01$ $11.2$ $13.13$ $11.69$
ru 12,112 $11.97$ $16.87$ $14.95$ $21$
zh 7,085 $3.58$ $6.39$ $4.4$ $5.67$
ar 2,283 Low $0.44$ $1.96$ $1.26$ $0.48$
cy 1,241 $0.64$ $1.24$ $1.21$ $0.72$
et 1,782 $0.49$ $1.46$ $0.61$ $0.93$
id 1,243 $0.31$ $0.67$ $1.27$ $0.79$
ja 1,119 $0.27$ $0.67$ $0.36$ $0.62$
lv 2,337 $1.22$ $1.65$ $2.63$ $2.06$
mn 2,067 $0.16$ $0.44$ $0.31$ $0.19$
nl 7,108 $2.14$ $6.95$ $10.75$ $8.46$
sl 1,843 $0.41$ $1.83$ $1.23$ $1.81$
sv 2,160 $0.9$ $1.28$ $3.4$ $1.76$
ta 1,358 $0.14$ $0.14$ $0.13$ $0.12$
tr 3,966 $1.67$ $3.5$ $3.04$ $3.14$
Table 8: Results of Table 1 in BLEU score for each language pair. The best result in each group is denoted by ’_’.