### v2

# Compact Speech Translation Models via Discrete Speech Units Pretraining

Abstract
We propose a pretraining method to use Self-Supervised Speech (SSS) model to creating more compact Speech-to-text Translation. In contrast to using the SSS model for initialization, our method is more suitable to memory constrained scenario such as on-device deployment. Our method is based on Discrete Speech Units (DSU) extracted from the SSS model. In the first step, our method pretrains two smaller encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Translation (DSU-to-Trl) data respectively. The DSU thus become the distillation inputs of the smaller models. Subsequently, the encoder from the Fbk-to-DSU model and the decoder from the DSU-to-Trl model are taken to initialise the compact model. Finally, the compact model is finetuned on the paired Fbk-Trl data. In addition to being compact, our method requires no transcripts, making it applicable to low-resource settings. It also avoids speech discretization in inference and is more robust to the DSU tokenization. Evaluation on CoVoST-2 (X-En) shows that our method has consistent improvement over the baseline in three metrics while being compact i.e., only half the SSS model size.
Compact Speech Translation Models via Discrete Speech Units Pretraining
Tsz Kin Lam  and Alexandra Birch  and Barry Haddow School of Informatics, University of Edinburgh {tlam, a.birch, bhaddow}@ed.ac.uk
1 Introduction
In Speech-to-text Translation (ST), using Self-Supervised Speech (SSS) models, such as wav2vec 2.0 and HuBERT Baevski et al. (2020); Hsu et al. (2021), as model initialization is now common to obtain the SOTA result Agarwal et al. (2023). Nevertheless, such model initialisation makes the ST model less memory-adaptive and could impose a large memory footprint. These factors hinders on-device deployment that is crucial for privacy and useful in the absence of internet connection.
How can we use the SSS model(s) to create a more compact ST model? When using the SSS model for initialization, the corresponding ST model uses the dense representations of the SSS model for its task. Alternatively, an informative proxy, which requires less memory to obtain, for the dense representation may make the ST model more compact.
Discrete Speech Units (DSU) extracted from the SSS model can be such a good proxy. DSU are K-Means clusters of speech representations from selected layers of the SSS model. It represents sequence of discrete tokens, which are easier to model within a text processing architecture Polyak et al. (2021); Chou et al. (2023). DSU sequences [^1] In this paper, DSU and DSU sequences are used interchangeably. When we need to focus on a few units of the sequence, we call them DSU tokens. are far smaller than the sequences of dense representations. Therefore, a straightforward method to distill the SSS models is to use DSU as speech inputs, aka the DSU-to-Translation (DSU-to-Trl) model. Although using DSU as inputs allows for transfer learning and a memory-adaptive model, using them at inference still requires storing and calling the quantization modules, i.e, the SSS model and the K-Means model.
We thus propose to use DSU for pretraining (PT) rather than as model input to make ST models more compact. Our method distils the SSS model by pretraining smaller models on the corresponding DSU. More specifically, our method firstly pretrains two smaller encoder-decoder models on 1) Filterbank-to-DSU (Fbk-to-DSU) and 2) DSU-to-Trl data respectively. The DSU thus become the distillation inputs of the smaller models. Subsequently, the encoder from the Fbk-to-DSU model and the decoder from the DSU-to-Trl model are taken to initialise the compact model. Finally, the compact model is finetuned on the paired Fbk-Trl data. Under this formulation, (1) we can use the SSS model to create a ST model that is adaptive to the memory footprint. (2) Our method requires no transcripts, unlike ASR-pretraining, making it applicable to low-resource settings. (3) Our method avoids using the quantization modules in inference. (4) Extensive results also show that our method is more robust to DSU tokenization than the DSU-to-Trl method.
We evaluate our method on CoVoST-2 Wang et al. (2021) X-En language directions ($21$ in total) using multilingual ST. By using a HuBERT-Base model to extract the DSU, our method shows strong and consistent improvements in three evaluation metrics with respect to a ST model that is trained from scratch. Our main contributions are:
  • We propose a pretraining method to distil the SSS model to creating a more compact ST model. Rather than competing with the SOTA ST models, adaptability to the memory footprint is our key focus.
  • Our method uses DSU for pretraining rather than as model inputs. This lowers the inference cost, especially for on-device purpose, by avoiding the quantization modules (storage and running).
  • We conduct extensive analysis to study the effect of DSU tokenization to both using DSU as model inputs and as pretraining. Our pretraining method is found to be more robust to different tokenizations.
2 Related Work
There are a number of related works that use DSU to enhance ST. Fang and Feng (2023) and Zhang et al. (2023b) use DSU to create more training data in a back-translation fashion. Chang et al. (2023) and Zhang et al. (2023b) explore the replacement of Filterbank by DSU as speech input. Furthermore, Yan et al. (2024) proposes a multi-tasking learning framework with hard parameter sharing, i.e., using a joint vocabulary for text tokens and DSU, to improve the speech-text modality gap. In contrast, we use DSU and its translation model for pretraining, resulting in a better Fbk-to-Trl model that has a shorter inference pipeline.
In the case of pretraining, Wu et al. (2023) use a single Speech-to-DSU model in pretraining for general speech-to-text purposes whereas we tailor the use for ST by using a pair of encoder-decoder models. Zhang et al. (2022b) also decompose ST into speech-to-unit and unit-to-text tasks. Their training is based on masked unit prediction, and it requires an extra unit-encoder module in inference. In contrast, we resort to supervised training on the DSU in acoustic pretraining and require no extra module in inference. More importantly, our goal is to make (multilingual) ST more compact, aiming also at low-resource settings where transcripts are not easily available, rather than learning a joint semantic space for both transcripts and audios.
3 Method
Figure 1: Illustration of the Fbk-to-DSU model. It is like an auto-encoding training process, but between a continuous format (log Mel Filterbank) and its discrete format (DSU) that is extracted from a HuBERT model.
Our method uses DSU in the form of pretraining to distil knowledge from the SSS (dense) representations to creating a more compact ST model.
In the first step, our method pretrains two smaller encoder-decoder models on 1) Fbk-to-DSU and 2) DSU-to-Trl data respectively. The Fbk-to-DSU model takes the log Mel Fbk as the encoder input and predicts the DSU sequence. The model is trained by an interpolation of Connectionist Temporal Classification (CTC, Graves et al. (2006)) loss that is applied to the last encoder layer and label-smoothed Cross-Entropy (CE) loss:
$\mathcal{L}^{\text{Fbk-to-DSU}}=(1-\lambda_{\alpha})\mathcal{L}_{\text{CE}}(\mathbf{U}|\mathbf{F})+\lambda_{\alpha}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{U}}|\mathbf{F})$ (1)
[…]
Next, we use the encoder of the Fbk-to-DSU model and the decoder (and its output layer) of the DSU-to-Trl model to initialise the compact model, followed by finetuning on the paired Fbk-Trl data using both CE and CTC loss Gaido et al. (2021); Zhang et al. (2023a) on the translations:
$\mathcal{L}^{\text{FT}}=(1-\lambda_{\beta})\mathcal{L}_{\text{CE}}(\mathbf{Y}|\mathbf{F})+\lambda_{\beta}\mathcal{L}_{\text{CTC}}(\mathbf{\tilde{Y}}|\mathbf{F})$ (2)
[…]
3.1 Tokenization of DSU in different models
Figure 2: Aligning the DSU tokenization of the Fbk-to-DSU, DSU-to-Trl and compact ST model.
The discrete nature of DSU makes the above training process similar to the transcripts-based pretraining. However, DSU is self-supervised, whereas transcripts require human annotations. DSU are also much longer and can be represented with various sets of symbols.
The length issue could be relieved by merging sequential repetitions Ao et al. (2022), e.g., ’#$1$$1$$1$$456$$456$$23$’ becomes ’#$1$$456$$23$’, where each DSU token is denoted by a #{integer}. Byte Pair Encoding (BPE) Sennrich et al. (2016) could be applied to reduce the DSU sequence length further, e.g., ’#$1$$456$$23$’ could be split into a single subword unit: ’#$1$#$456$#$23$’.
Since both Fbk-to-DSU and DSU-to-Trl models map to different targets, and DSU can be represented with various set of symbols, we align the tokenizations (or called vocabularies [^2] We use vocabulary and tokenization interchangeably, since we did not apply subword regularisation.) of the two models. Figure 2 provides an illustration. The vocabulary of the Fbk-to-DSU model (Vocabulary A) is identical to the source vocabulary of the DSU-to-Trl model (their weights are not shared since these two models are trained independently), whereas the target vocabulary (Vocabulary B) of the DSU-to-Trl model is identical to the target vocabulary of the final compact model (their weights are shared during initialisation). The DSU-to-Trl model is similar to a text translation model, so we also experiment of using separate vocabularies or a joint vocabulary. If a joint vocabulary of English subword units and DSU (BPE or not) is used, all the three models would have the same vocabulary, and the weights of the source and target vocabularies of the DSU-to-Trl model are also tied.
4 Experiments
4.1 Data Preprocessing
We follow standard practices to preprocess the CoVoST-2 X-En data. For speech inputs using $80$-D log Mel Fbk, we computed the features for every $10$ms with a $25$ms window and then normalized them using its mean and variance computed over each channel. We use the BPE implementation from sentencepiece Kudo and Richardson (2018) and obtain vocabulary of size $8$K on the English target, $16$K on the (non-English) transcripts and $32$K on the DSU, unless otherwise specified.
We use HuBERT-Base [^3] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/gslm/speech2unit model to extract the DSU by first downsampling the CoVoST-2 audio to $16$KHz. Each audio data utterance is then converted into the DSU, i.e., the clustering indexes, by applying K-Means clustering (K=$1,000$; MiniBatchKMeans from sklearn) on its HuBERT representation from the $6$th layer Lakhotia et al. (2021). To train the K-Means model, we divide the $21$ language pairs into three groups: 1) {ar, cy, et, id, ja, lv, mn, sl, sv, ta, tr}, 2) {nl, pt, ru, zh} and 3) {ca, de, es, fa, fr, it}. We then sample $1$K instances for each language pair in group 1), which becomes $3$K in group 2) and $12.5$K in group 3), to create a multilingual training dataset of $98$K instances for the K-Means model.
4.1.1 On the choice of using HuBERT-Base
Given the rapid advance in the SSS models, there are many alternatives, such as XLS-R Babu et al. (2021) and Wavlm Chen et al. (2022), for extracting the DSU for our method. These models are larger in scale and could be multilingual, thus providing DSU of higher qualities. The improvement of our method by using DSU from the HuBERT-Base would probably be a lower-bound, considering its relatively poor qualities to the bigger models. Since our goal is about compactness via DSU pretraining rather than comparing the DSU qualities across the SSS models, we took a simple HuBERT-Base model to illustrate the idea. Pretraining only on English audio data could also suggest hints on whether the DSU and our method could be generalised to languages that are unseen to the SSS models.
4.2 Model Configuration
All models are based on Transformer Vaswani et al. (2017) with implementations from fairseq Ott et al. (2019); Wang et al. (2020). In the Fbk-to-Token (i.e. transcriptions, DSU, or translations) models, the encoder has convolutional layers to downsample the Fbk by a factor of $4$. There are $12$-$6$ layers in the transformer encoder-decoder, whereas the embedding and feed-forward network (FFN) dimensions are $256$ and $4,096$ respectively, unless otherwise specified. It is worth noting that: (1) The Fbk-to-DSU model is not trained on the translations, so it is not directly comparable to the ST models. Its effect on ST lies on its pretrained encoder (Table 2). (2) The DSU-to-Trl model is a ST model which decoder can be used for initialization.
Scratch is a ST model trained on the paired speech-translation data without pretraining.
ASR Pretraining refers to a ST model whose encoder is initialized by a speech recognition task with CTC regularisation on the transcripts.
DSU-to-Trl follows the Transformer used in text translation. We use $6$-$6$ layers in the encoder-decoder which the dimension of embedding and FFN is $256$ and $2,048$ respectively. In addition, we use "pre" layer-normalization Nguyen and Salazar (2019). Despite its smaller model size, its inference requires the quantization modules.
Hu-Transformer uses the entire HuBERT as the speech encoder initialization Fang and Feng (2023). For comparison to our DSU-Adapter, its subsequent encoder-decoder also has $1$-$6$ layers.
DSU-Adapter is our proposed method. To better align the two pre-trained components, we also experiment with adding an extra encoder layer as a simple adapter layer after the pre-trained encoder. Because of the small model size, all model parameters are trainable. Since its decoder is initialized by the DSU-to-Trl method, its decoder FFN dimension is $2,048$.
Enc-Init is a ST model that has its encoder initialized by the Fbk-to-DSU encoder. EncDec-Init is a DSU-Adapter model without the adapter layer.
4.3 Training and Inference
It is worth noting that we do not use extra audio data, e.g., Libri-Light Kahn et al. (2020) in our (pretraining) experiments. Furthermore, we apply the following conditions in (pre-)training:
  • We skip training data that are longer than $30$ seconds (audio) or $1,024$ target tokens.
  • We apply SpecAugment Park et al. (2019) with parameters: {$F=30,T=40,m_{F}=2,m_{T}=2$} on Filterbank inputs.
  • We share the embedding weights when using a joint vocabulary in the DSU-to-Trl model.
  • We set $\lambda_{\alpha}$ and $\lambda_{\beta}$ in CTC to $0.3$ and the smoothing parameter to $0.1$
  • We initialize the encoder (decoder) with the last (best) checkpoint from the PT model.
  • We use Adam optimizer with inverse square root scheduler for all model training.
  • In all Fbk-to-Token models, the effective mini-batch size, warm-up steps, peak learning rate and training steps are $32$K frames, $25$K, $2$e33-3- 3 and $60$K steps respectively.
  • Similarly, in all DSU-to-Trl models, we use $80$K tokens, $10$K, $5$e44-4- 4 and $50$K steps.
  • Similarly, in Hu-Transformer, we use $4$M frames, $4$K, $1$e44-4- 4 and $300$K steps.
In inference, we average the last 5 checkpoints and use beam size of $5$ in generation. All experiments are run on Nvidia A100 GPUs. It takes about $1$ day for $2$ A100 ($40$GB) GPUs to complete an experiment that uses Filterbank as speech inputs.
AST model (#Params) BLEU chrF COMET-22-DA
High Mid Low All High Mid Low All High Mid Low All
Scratch (52M) $19.4$ $7.91$ $0.73$ $5.99$ $43.6$ $27.2$ $14.6$ $23.1$ $0.605$ $0.498$ $0.433$ $0.481$
ASR-Pretraining (52M) $26.5$ $12.2$ $1.82$ $9.00$ $51.9$ $32.8$ $16.4$ $27.1$ $0.680$ $0.537$ $0.443$ $0.511$
Hu-Transformer (113M) $24.3$ $11.4$ $2.18$ $8.60$ $49.9$ $31.9$ $17.0$ $26.8$ $0.650$ $0.522$ $0.439$ $0.499$
DSU-Adapter (48M) $26.5$ $12.9$ $1.76$ $9.13$ $52.1$ $33.9$ $16.5$ $27.4$ $0.681$ $0.548$ $0.442$ $0.513$
Table 1: Results in BLEU, chrF and COMET-22-DA on the test set of CoVoST-2 (X-En) by resource group. In all metrics, DSU-Adapter is much better than Hu-Transformer, which is $2.3$ times larger, in both "High" and "Mid" groups. DSU-Adapter, which does not requires transcripts in training, is also on a par with ASR-Pretraining. The best result in each group is denoted by ’_’.
5 Results and Analysis
Before discussing the results, it is worth noting that (1) Hu-Transformer is not memory-adaptive, and (2) ASR-Pretraining requires transcripts, unlike DSU which is self-supervised. Both methods are introduced for reference purposes of if such resources are available.
5.1 Improvement brought by DSU-Adapter
We divide the 21 language pairs by resource level into: 1) "High": {ca, de, es, fr}, 2) "Mid": {fa, it, pt, ru and zh}, 3) "Low": {ar, cy, et, id, ja, lv, mn, nl, sl, sv, ta, tr} and 4) "All": the 21 languages pairs. We report the average BLEU444nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.3.1 and chrF555nrefs:1—case:mixed—eff:yes—nc:6—nw:0—space:no—version:2.3.1 over the test sets of each group using sacrebleu Post (2018). In addition, we also provide the result in WMT22-COMET-DA Rei et al. (2022), which the source inputs are the gold-reference transcripts.
Table 1 compares our DSU-Adapter and the baselines. Our DSU-Adapter is $3$ BLEU (in the group "All") higher than the Scratch model. This shows that our proposed method of using DSU-pretraining can strengthen direct end-to-end ST without requiring transcripts and remain flexible in memory footprint (smaller in size than the HuBERT model). Furthermore, it is better than Hu-Transformer in spite of having half the parameters. For "Mid" and "High", the improvement in BLEU is $1.49$ and $2.23$ points respectively, but it falls short by $0.42$ points for "Low". We also compare to ASR pretraining, which is not always applicable, e.g., in low-resource setting or perhaps even in an unwritten language Zhang et al. (2022a). Surprisingly, our adapter is on a par with it, and its BLEU is $0.13$ points better. The result remains consistent when it is measured in chrF and COMET.
5.1.1 Language-specific performance
Figure 3 shows the performance on each language pair in BLEU, chrF and COMET-22-DA. Our DSU-Adapter (in green triangles) show consistent improvement over the Scratch model (in blue circles) in all language pairs. Such improvement is rather surprising since HuBERT-Base was trained solely on English audio data. We hypothesized that the cross-lingual improvement is related to HuBERT’s ability to capture language independent features, e.g. phonetic properties Pasad et al. (2023).
Compared with Hu-Transformer, DSU-Adapter maintains an evident improvement over most language pairs in both "High" and "Mid" groups. Exceptions are in "fa" and "pt", but the lags are almost negligible. In group "Low", Hu-Transformer is slightly better, especially in "nl" and "sv" pairs. However, most translation in this group is barely around $2$ BLEU, and the lags are small.
In most language pairs, DSU-Adapter performs similarly to ASR-pretraining (in red diamonds), except translating from "ru" audios. The improvement in this "ru-en" pair makes DSU-Adapter to have an evident advantage of $0.7$ BLEU in the group "Mid".
Has BPE $|\mathcal{V}|$ DSU Length DSU-to-Trl Enc-Init EncDec-Init DSU-Adapter
on DSU? Length Ratio ($20$M to $27$M) ($52$M to $70$M) ($46$M to $64$M) ($48$M to $67$M)
No $1$K-$8$K $176$ $12.9$ $6.73$ $7.70$ $7.87$ $8.54$
$1$K-$16$K " $14.1$ $6.36$ $7.50$ $8.00$ $8.43$
$1$K-$32$K " $14.9$ $6.30$ $7.23$ $7.65$ $7.94$
$8$K " $12.7$ $6.88$ $7.64$ $8.05$ $8.26$
$16$K " $14.0$ $6.33$ $7.41$ $7.91$ $8.17$
$32$K " $14.9$ $6.26$ $6.66$ $7.41$ $7.68$
Yes $1$K-$8$K $221$ $16.3$ $4.52$ $8.23$ $8.44$ $8.61$
$16$K-$8$K $129$ $9.5$ $5.06$ $8.51$ $8.76$ $8.95$
$32$K-$8$K $115$ $8.5$ $4.43$ $8.67$ $9.02$ $9.13$
$8$K $150$ $7.6$ $7.02$ $8.33$ $8.51$ $8.82$
$16$K $133$ $7.7$ $6.50$ $8.57$ $8.61$ $8.93$
$32$K $118$ $7.8$ $5.07$ $8.30$ $8.44$ $8.70$
Table 2: (DSU) tokenization effect on 4 ST methods. Each ST model’s performance on the CoVoST-2 test set is measured by BLEU on group "All". All 4 methods could perform better than the Scratch model of $5.99$ BLEU as shown on Table 1. In general, darker (brighter) cells refer to weaker (stronger) models. The best two models apply both BPE on the DSU and separate vocabularies in PT (cells in yellow).
5.2 Tokenization effect to the DSU-to-Trl method and the DSU-Adapter method
In this section, we investigate how tokenization, including BPE, affects the DSU-to-Trl method and the DSU-Adapter method. We are particularly interested in their robustness toward the tokenization, especially using BPE on the DSU, since tuning the quantization process and retraining the subsequent models is computationally expensive.
In Table 2, the 1st column "Has BPE on DSU?" indicates if BPE is applied on the DSU. If "Yes", multiple DSU could be merged into one subword unit, e.g., ’#1 #456 #23 #999’ could be merged into ’#1#456#23#999’. The 2nd column "$|\mathcal{V}|$" shows the vocabulary configuration: its size, and if the model has a joint vocabulary. For example, "$1$K-$8$K" means that we use a vocabulary of size $1$K for DSU and a second vocabulary of size $8$K for English so that the DSU-to-Trl model would have separate vocabularies for the source (DSU) and target (English) sides. All results are in BLEU averaged over all language pairs, i.e., group "All".
5.2.1 DSU-to-Trl: robust to tokenization?
When BPE is not applied on the DSU, those 6 DSU-to-Trl models have $6.48$$\pm$ $0.26$ BLEU. Despite having smaller model size (<$30$M), they are better than the Scratch model of $5.99$ BLEU.
When BPE is applied, the sequence length of DSU (DSU Length) could be shortened, which could in turn improve the performance, e.g. the best DSU-to-Trl model happens at configuration "$8$K" with $7.02$ BLEU. However, the DSU-to-Trl method is quite unstable to the use of BPE, as reflected by the $5.12$$\pm$ $0.83$ BLEU in the other 5 configurations. The correlation between the DSU sequence length, the source-target length ratio, and the ST performance is also not straightforward. For an example, the "$32$K" model (DSU length of $118$) is about $2$ BLEU behind to the "$8$K" model (DSU length of $150$). Therefore, applying BPE on the DSU for length reduction should remain cautious.
5.2.2 The DSU-Adapter is more robust
Unlike DSU-to-Trl method, DSU-Adapter benefits more when BPE is applied to the DSU. Our proposed method has $8.86$$\pm$ $0.19$ BLEU (over the 6 corresponding configurations), as opposed to $8.17$$\pm$ $0.32$ BLEU when BPE is not applied. This observation is opposed to the DSU-to-Trl method which only scores $5.54$$\pm$ $1.07$ BLEU (with also larger variance) when BPE is applied on the DSU but $6.48$$\pm$ $0.26$ when BPE is not used. The improved mean score and its smaller variance suggests that the DSU-Adapter method is more (DSU) tokenization robust. We see this as a benefit of introducing the DSU, i.e., the SSS model knowledge, via PT rather than as model inputs.
On top of applying BPE on the DSU, using separate vocabularies in PT is preferred (the two yellow cells on Table 2) since it performs slightly better, and the DSU, which are not needed in the ST output, would not occupy the target vocabulary.
5.2.3 Ablation: initialisation in DSU-Adapter
Having similar model sizes, e.g. about $50$M parameters (Table 2), DSU-Adapter is better than both EncDec-Init and Enc-Init methods. The translation performance in BLEU (averaged over the 12 vocabularies) is $8.51$$\pm$ $0.44$, $8.22$$\pm$ $0.51$, and $7.99$$\pm$ $0.61$ respectively. Encoder-initialization seems more crucial than decoder-initialization, as reflected by the fact that the best DSU-Adapter model comes from a combination with the weakest DSU-to-Trl model of $4.43$ BLEU.
5.3 Is CTC applicable also to DSU?
Has CTC in High Mid Low All
DSU PT? ST FT?
No No $25.81$ $9.91$ $1.46$ $8.14$
No Yes $26.10$ $11.35$ $1.69$ $8.71$
Yes No $25.94$ $10.82$ $1.51$ $8.44$
Yes Yes $26.12$ $11.53$ $1.73$ $8.74$
Table 3: Effect of CTC on Fbk-to-DSU PT and/or ST FT to the DSU-Adapter method. All results are in BLEU and the best in each group is denoted by ’_’.
Similar to ST methods that use pretrained components, our method could be limited by the pretraining modality gap Liu et al. (2020); Le et al. (2023). Motivated by prior works, we investigate mitigating it with CTC. A crucial difference to the prior works is that our method uses DSU for pre-training rather than transcripts.
We thus study applying CTC in our method at different training stages. Owing to the large number of vocabulary configurations on Table 2, we only experiment with: 1) "No-BPE $1$K-$8$K", 2) "BPE $8$K", 3) "BPE $32$K" and 4) "BPE $32$K-$8$K". In each training stage, we report the effect of CTC to the ST performance (per resource group) by averaging the BLEU of these $4$ configurations.
Table 3 presents the analysis of applying CTC on our DSU-Adapter method. The training condition "Has CTC in DSU PT" refers to the case of applying CTC on the discrete speech units in Fbk-to-DSU pretraining, whereas "Has CTC in ST FT" refers to the case of applying CTC on the translations in ST finetuning, i.e., on the paired Fbk-Trl data. Our result shows that CTC helps on either stage, but the gain is $0.27$ BLEU more in ST finetuning. Using them jointly still helps, but the marginal gain is barely $0.03$ BLEU.
6 Limitations and future works
In the previous sections, we discuss the noticeable benefits of our DSU-pretraining method in creating a more compact ST model. In spite of this, there are several factors that are not thoroughly explored and could improve the model performance further:
K-Means clustering
We did not inspect the clustering size (fixed to $1,000$) and the number of training instances (only fixed to $98,000$) used in training the K-Means clustering model. Apart from tuning its hyper-parameters, using other techniques, such as residual vector quantisation Zeghidour et al. (2021); Défossez et al. (2022) and multiple codebooks Guo et al. (2023), might bring better improvement.
Other acoustic encoders
We did not experiment other acoustic encoders, such as conformer Gulati et al. (2020); Papi et al. (2023) and E-Branchformer Peng et al. (2023). This stronger encoders should provide further gains for our method since they also enjoy the benefit of pretraining.
A stronger pretrained decoder
Apart from strengthening the encoder, the DSU-to-Trl model and hence its decoder (used in initialisation) could also be improved, e.g. via back-translation, up-sampling the textual sequence Yan et al. (2024) and pretraining with more text data, while maintaining the small decoder size.
Further analyses
In addition to improving our pretraining method for better model compactness, there are other related research directions worth further analyzing. One direction would be how, in terms of acoustic pretraining, DSU compared with transcripts (if available in that language) over different data scales. Another interesting research direction would be the comparison and analysis of using DSU or dense features in a large pretrained model setting, such as Whisper Radford et al. (2023) and Large Language Models.
7 Conclusion
In this paper, we consider a memory-constrained setting for ST. Our proposed method uses DSU in the form of pretraining to distil the knowledge from the Self-Supervised Speech model to creating more compact Speech-to-text Translation. Our compact model, i.e., the DSU-Adapter, shows strong and consistent improvements in three evaluation metrics over the baselines. In contrast to using DSU as model inputs, our method does not require quantization modules in inference and shows stronger robustness to the DSU tokenization. Finally, our method requires no transcripts, making it also suitable for low-resource setting.
Figure 3: Results in BLEU, chrF and COMET-22-DA on each language pair of CoVoST-2 (X-En).
Acknowledgements
This work was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10039436: UTTER). The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.